Consider the heat released from the rubbing alcohol bottle is 50,000J as its temperature changed from 25°C to 86°C. Calculate the mass of rubbing alcohol used into the bottle if the specific heat of rubbing alcohol is 2.3J/gC.
1
Expert's answer
2020-11-09T13:58:39-0500

Q=m*C*(t2-t1);

50000 = m *2.3*(86-25)

m=50000/(2.3*61) = 356.38 g